Answer the following questions about marginal propensity to consume and the multiplier. First provide the correct equation and then show your work to arrive at the answer:


What is the marginal propensity to consume when consumption changes from 7 to 6 and disposable income changes from 5 to 3?

If disposable personal income is 10 and consumption is 12, what is personal savings? What does this mean?

What is the multiplier when the change in equilibrium level of real GDP in the aggregate expenditures model is 9, and change in autonomous aggregate expenditures is 3?

What is the multiplier when the marginal propensity to save is 1/3?

What would happen to the marginal propensity to save when a tax cut was enacted causing the multiplier to change to 5?


1
Expert's answer
2021-12-01T10:09:01-0500

If the change in consumption is from 7 to 6, delta C is 1. If the change in income is from 5 to 3, delta Y is 2. Delta C / Delta Y = ½. The marginal propensity to consume is therefore 0.5.


Savings = Earnings - Consumption. = 10 - 12. = -2 units. This means that savings cash is reducing by 2 units each time consumption is done.


The Multiplier equation is as follows: Multiplier = Change in Equilibrium Level of real GDP/ Autonomous Aggregate Expenditure = 9/3 = 3